please I need help with this question;

7. Titration of Acetic Acid: using the equation [H^+] = (ka [HA])^1/2 for weak acids, calculate the theoretical starting pH of the titration based on the concentration of acetic acid notated on the bottle and the final concentration of the solution after preparing it for the reaction. calculate the percent error relative to the pH that was recorded in the lab.

(the concentration of the acetic acid is 0.10M and the pH recorded in the lab is 3.7)
Answer #1

Ka of acetic acid, CH3COOH is Ka = 1.8*10-5.

Write down the ionization of CH3COOH as below.

CH3COOH (aq) <======> H+ (aq) + CH3COO- (aq)

The acid ionization constant is expressed as

Ka = [H+][CH3COO-]/[CH3COOH]

======> 1.8*10-5 = (x)(x)/(0.1 – x)

where x is the equilibrium concentration of H+ = equilibrium concentration of CH3COO- due to 1:1 nature of ionization of acetic acid.

Assume x << 0.1 M (since Ka is small).

Therefore,

1.8*10-5 = x2/(0.1)

=====> x2 = 1.8*10-5*0.1

=====> x2 = 1.8*10-6

=====> x = 1.3416*10-3

Therefore,

[H+] = 1.3416*10-3 M

and

pH = -log [H+]

= -log (1.3416*10-3 M)

= 2.8723 ≈ 2.87

The initial pH of the acetic acid solution is 2.87 (ans).